HOW TO SOLVE MENTALY THIS IN 5 SECONDS 68*69
Answer by ewatrrr(24785)


Hi
HOW TO SOLVE MENTALY THIS IN 5 SECONDS 68*69
the 'Quick' Idea is based on 68*69 = (70-2)(70-1) = 4900-3*70 +2 ..4690 + 2
4692